OpenStudy (anonymous):

Really need someone to check my answer. The variable Z is inversely proportional to X. When X is 6, Z has the value 2. What is the value of Z when X=13 Round to nearest thousandth if needed. My answer was 4 333

OpenStudy (whpalmer4):

If Z is inversely proportional to X, then we can write \[Z = \frac{k}{X}\]where \(k\) is our proportionality constant. We know a pair of values for \(X\) and \(Z\), so we can plug them into the equation and solve for \(k\), then plug the second value of \(X\) in and solve for \(Z\). \[Z =\frac{k}{X}\]\[2=\frac{k}{6}\]\[k=2*6=12\]\[Z = \frac{12}{X} = \frac{12}{13}\approx\]

OpenStudy (mathmale):

Fab: You are told that when x is 6, z has the value 2. Now that you have \[z=\frac{ 12 }{ x},\]

OpenStudy (mathmale):

check this equation by substituting the x and y values given. Is the equation true? If yes, the equation is correct. So, you see, you can check your work here yourself. Next, substitute x=13 and calculate z.